OpenStudy (anonymous):

How do I reverse the order of integration of the following double integral:

OpenStudy (anonymous):

bring it on. you should always draw the area btw

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{x} 1-x^2 dydx\] I guessed it would simply be to change the bound to: \[\int\limits_{0}^{1}\int\limits_{0}^{y}dxdy\]

OpenStudy (anonymous):

(same function in the second integral)

OpenStudy (anonymous):

The region is basically the triangle bounded by 0<x<1 and the straight line y=x

OpenStudy (anonymous):

that is right. the way i do it is by drawing it:  then i look at it sideways y is bounded between x = 0 and x = 1 : 0 < y < 1 and y is bounded between y = x and y = 1: x < y < 1

OpenStudy (anonymous):

lol i just wrote same bounds twice

OpenStudy (anonymous):

second on is: x between y and 1 y < x < 1 sorry about that

OpenStudy (anonymous):

I messed up the bounds, but I understand now. Thank you!